OpenStudy (anonymous):

Please help quickly!!! Find the slope of the line containing the following two points (5/8,-2) and (-1/4,-1/4)

OpenStudy (lgbasallote):

a line with two given points has slope (y2-y1)/(x2-x1) where x1 and y1 are the first x and y coordinates respectively; x2 and y2 are the second x and y coordinates respectively

OpenStudy (anonymous):

So the slope is 2/2?

OpenStudy (anonymous):

no.

OpenStudy (lgbasallote):

seems like a fraction is the answer...

OpenStudy (anonymous):

Rise over run..

OpenStudy (lgbasallote):

-1/4 - (-2) ---------- -1/4 - 5/8 -1/4 + 8/4 ---------- -2/8 - 5/8

OpenStudy (anonymous):

The slope would be -7/8?

OpenStudy (anonymous):

Find the rise fist. This is y2-y1. Find the run second. This is x2-x1. Find rise divided by run.

OpenStudy (lgbasallote):

@SmoothMath your y1 should be -2...not -5/8 ^_^

OpenStudy (anonymous):

Oh haha yes. I flipflopped rise and run. Let me try that again.

OpenStudy (anonymous):

(5/8,-2) and (-1/4,-1/4) Rise: x2-x1 = (-1/4) -(-2) = (-1/4) - (-8/4) = 7/4 Run: y2-y1 = (-1/4) -(5/8) = (-2/8)-(5/8) = -7/8 Haha how embarassing.

OpenStudy (anonymous):

Deleting incorrect posts, just to avoid confusion.